1. You are buying a new television. From past experience, you estimate future repair costs as: First Year: $5 Second year: $15 Third year: $25 Fourth year: $35

A.The dealer offers to sell you a four-year repair contract for $60. You require at least a 6% interest rate on your investments. What is the Annual uniform series ($) of this four-year repair costs (with no contract)?

B. The dealer offers to sell you a four-year repair contract for $60. You require at least a 6% interest rate on your investments. What is the future value ($) of this four-year repair costs (with no contract)?
Answer #1

Answer 1(A):

Given that:

Future repair costs as:

First Year: $5

Second year: $15

Third year: $25

Fourth year: $35

Interest rate = 6%

With no repair contract:

Present value of Repair cost = 5/1.06 + 15 / 1.06^2 + 25/1.06^3 + 35/1.06^4

= $66.78

Annual uniform series ($) = Present value / PV factor for annuity

= 66.78 / ((1 - 1/1.06^4)/6%)

= $19.27

Annual uniform series ($) of this four-year repair costs (with no contract) = $19.27

Answer 1(B):

Future value ($) = 5*1.06^3 + 15 *1.06^2 + 25*1.06 + 35

= $84.31

Future value ($) of this four-year repair costs (with no contract) = $84.31